How to Use the gradient from an equation of Average mass vs Average radius cubed r3 to calculate the density
Answer #1

Cube each side of the equation, giving you:

R3 = r03A

You then need to calculate R3 and plot R3 against A. That will give you a straight line graph with a gradient of r03.Calculate the gradient and value of r0(cube root of the gradient).

For density calculation:

Assuming that-

•the nucleus is spherical OR

•all nuclei have the same density OR

•that total mass is equal to the mass of constituent single nuclei (ignoring the mass difference) OR

•ignoring the gaps between nucleons

Proceeding using the density formula (i.e. rho=M/V)

where rho=density, V=4/3 (π R3)- volume of sphere, , M= mass of a nucleon.

Putting the value of V, we get rho= 3/4(M/πR3)

If A=1 then R=r0(here to use the gradient value)

Putting the value of R in the given equation and calculate the density.